OpenStudy (anonymous):

. Divide (x^2+6x+9)÷(x+3) A. 0 B. 1 C. (x + 3)^2 D. (x + 3)

OpenStudy (anonymous):

D. (x + 3) is the correct answer. @LaBlueEyes2013

OpenStudy (anonymous):

\[ (x^2+6x+9)÷(x+3) = \frac{ (x^2+6x+9)}{(x+3)}= \frac{ (x^2+6x+3^2)}{(x+3)}\] \[= \frac{ (x+3)^2}{(x+3)} =(x+3)\] @@LaBlueEyes2013